# Physical properties of strong 1 ¡ z ¡ 3 Balmer and Paschen lines emitters observed with JWST

### v1

@@ 3.2 Analysis of observed line ratios
We want to ascertain if some objects in our sample have hidden attenuation before going further into our analysis. Indeed, a galaxy with a high Pa/H ratio compared to a Balmer decrement close to 2.86 would indicate a hidden and highly attenuated component. To do this, we compare the observed line fluxes ratios to the theoretical ones. However, this comparison comes with a few caveats, the first one being slit losses. Following Napolitano et al. (2024) we assume that the slit losses affect all lines equally. The second issue comes from the underlying absorption lines of the stellar continuum, especially for H, H and H which have an absorption line EW 5Å(see Table 3). We correct the fluxes for these absorptions by computing the absorbed flux from the best model of CIGALE (Sect.3.3).
@@
We find that our objects encounter, at most, a very moderate hidden attenuation, a result comparable to what Reddy et al. (2023) found for a sample of similar redshift CEERS sources. Taking into account the uncertainties on the flux measurements, our galaxies have either a very low attenuation or are compatible with the attenuation model presented in Sect 3.1. From Fig 3, we conclude that the modified CF00 recipe is able to reproduce the attenuation processes at work in obscured galaxies.
@@ 3.3 Absorption line EWs
One common difficulty of measuring hydrogen emission lines EWs is the presence of an underlying absorption line. In our fitting method, presented in Sect.4.3, we reproduce the EW measurements carried out on observed spectra (Sect.2.1), avoiding any correction for an underlying absorption either on observed or model spectra.
Absorption lines are part of the stellar emission templates used to fit the galaxies and in this section we make use of the best fit model to compute seven hydrogen absorption lines EWs for each galaxy of our sample. We use the binary models of BPASS as even the resolution of the high resolution models from Bruzual & Charlot (2003) (BC03) is not enough to accurately model these absorption lines. We show an example of the best model of ID 5430 in Fig 4 and the values of the EWs are reported in Table 3.
@@
The main driver of the strength of absorption line EWs is likely to be the SFH of the galaxy. We find a continuous decrease of the strength of the absorption lines with increasing wavelengths, from Pa to H. We explain this result by a larger contribution of old stars to the stellar emission at longer wavelength, reducing the impact of absorption lines from late B and A stars on the stellar continuum.
@@ 4 Spectral energy distribution fitting
To fit the SEDs from the combination of broadband fluxes and EWs, we use the modelling software CIGALE. For a complete description of CIGALE and its functionalities; see Boquien et al. (2019). The models are built by successively calling modules, each corresponding to a single physical component or process. CIGALE combines a stellar SED built with an SFH, stellar populations models and nebular emission with dust absorption and emission components. The energy balance between stellar and nebular dust absorption and dust re-emission is conserved. Energetic photons produced by massive stars ionise the surrounding gas which re–emits the energy in the form of a series of emission lines and a nebular continuum.
The quality of the fit is assessed by the value of the and the values of the physical parameters and their corresponding uncertainties are estimated as the likelihood-weighted means and standard deviations. Below, we briefly present the modules we use and our input parameter values which are summarised in Table 4.
Our SED fitting combines photometric fluxes and EWs as spectroscopic data. When combining integrated photometric data and EWs measured on the NIRSpec spectra, we assume that these EWs are representative of the whole galaxies.
Along with the SED fitting process with CIGALE it is possible to build and analyse a mock catalogue related to the observed dataset(Buat et al., 2014; Ciesla et al., 2015). CIGALE uses the best-fit model of each of the objects obtained with the SED-fitting procedure. The flux densities and EWs of the mock SEDs are then computed by randomly picking a flux or EW value from the normal distribution generated using the best model value and the error of the input data as standard deviation. CIGALE is then run on this artificial catalogue with the same configuration as for the first run in order to compare the exact values of the physical parameters corresponding to the artificial SEDs to the parameters estimated by the code with the probability density function of each parameter. Below, we use this comparison to check the robustness of the output SED fitting parameters of interest for our study and our ability to constrain them.
@@ 4.1 Stellar population models and star formation history
We consider two models of single stellar populations: the widely used BC03 models and the BPASS models (version 2.2) which account for binary stars (Eldridge et al., 2017; Stanway & Eldridge, 2018). The latter has been shown to be very efficient when fitting UV and optical photometry with spectroscopic data (e.g. Larson et al., 2023; Bolamperti et al., 2023; Rezaee et al., 2023; Tacchella et al., 2023; Lecroq et al., 2024) as it was built with a focus on best reproducing the emission from young and massive stars (e.g. Eldridge et al., 2017; Stanway & Eldridge, 2018; Reddy et al., 2022; Marchant & Bodensteiner, 2023). The BPASS binary models have known issues when fitting stellar populations older than a few gigayears (Eldridge et al., 2017; Stanway & Eldridge, 2018; Han & Han, 2019), the first one being that models remain bluer compared to single star models. Two reasons are given by Eldridge et al. (2017) to explain this effect: the fixed binary fraction regardless of the mass of the star and the width of the time bins at late ages combined with a relatively sparse sampling of stellar masses. Binary models also take a longer time to reach the quiescence (Eldridge et al., 2017). This leads to an underestimation of the number of old and small binary stars.
In both cases, we choose the Initial Mass Function of Chabrier (2003) and set the metallicity to the solar value: .
@@ 4.2 Dust attenuation and emission
To account for the dust attenuation, we apply the model presented in Sect. 3.1 which corresponds to the modified_CF00 module in CIGALE. In this module, , , and (Eq. 1, 2, and 4) are free parameters. We sample the parameter space of , and very thoroughly (with 0.1 step). ranges from -0.4 to -1.5 to represent the broad distributions of values found from both observations (e.g. Lo Faro et al., 2017; Buat et al., 2019; Pantoni et al., 2021; Seillé et al., 2022) and simulations (e.g. Chevallard et al., 2013; Roebuck et al., 2019; Trayford et al., 2020). The tests we performed on mock catalogues show that is not constrained. We thus introduce only three values for : the widely used value of -1.3 (Charlot & Fall, 2000; da Cunha et al., 2008; Chevallard et al., 2013), -0.7 as proposed in the original CF00 recipe (Charlot & Fall, 2000) and an in-between value of -1.0.
To model the dust re-emission of the five sources that have far-IR data, we use the models from Draine & Li (2007), updated in Draine et al. (2014) which can be found in the dl2014 module in CIGALE. The model is built as follows: a fraction (1 - of the dust mass is heated by a starlight intensity , while the remaining fraction is exposed to starlight with intensities ¡ U ¡ , with a power-law distribution . The polycyclic aromatic hydrocarbon (PAH) abundance, , is also a free parameter in the models. As suggested by Draine et al. (2014), is fixed to and the slope of the power law, , to 2.
@@ 4.3 Nebular emission and EWs
The nebular emission in CIGALE is computed with CLOUDY 17.01 (Ferland et al., 1998, 2017) with the configuration described in Villa-Vélez et al. (2021) and Theulé et al. (2024). 129 nebular emission lines intensities are calculated and re-scaled with the number of Lyman continuum photons from the stellar emission of the galaxy. The ionising spectrum is modelled as a constant burst of star formation of 10 Myrs (Boquien et al., 2019; Villa-Vélez et al., 2021) computed with either BC03 or BPASS models.
We model the Paschen and Balmer lines EWs 444The EWs computation used in this work is not included in the current released version of CIGALE, 2022.1. The newest version of the EWs computation is available on request. so as to reproduce the measurements described in Sect 2.1. For each emission line of each galaxy, the input wavelengths (a six values ”array”) defined with LiMe are used to delimit the line region and two nearby featureless continua on each side of the line. The code then performs a linear fit of the overall continuum which will be subtracted from the line. LiMe deblends the H and [NII] doublet of our galaxy sample (see Fig 1), and these lines are also considered separately in CIGALE. Finally, as pointed out in Sect 3.2, these EWs are not corrected for underlying absorption as they are computed and thus affected in the same way as the EWs measured with LiMe.
@@ 5.1 Fit of the spectro-photometric data of BC03 vs BPASS
We adopt the limit of ¿ 6 to put a strong preference on a model, a positive preference being defined as 2¡ ¡ 6 (Salmon et al., 2016). The distribution is shown in Fig 6. Eleven objects present a strong preference (and one, ID 10293, a positive preference) for the BC03 models while only one galaxy (ID 23542) is found with a strong preference for the BPASS models. For this object, the photometric data is better fitted with the BC03 models (see the SEDs in Appendix A), but the EWs are better fitted with the BPASS models. The uncertainties on the measured EWs are very small for this galaxy ( times lower than for the other sources), which leads to a higher impact of spectroscopic data on the calculation of the value.
We present a more detailed comparison of the quality of the fit of the photometric data with BPASS and BC03 models in Fig 7. For each filter and both models, we compute the ratio between the bayesian estimate of the flux given by CIGALE over the measured flux. While the CFHT and HST fluxes do not show a significant difference of quality between both models, there is a clear trend for the four bluest NIRCam filters to underestimate the model fluxes with respect to the observed values when the binary BPASS models are used. The CFHT and HST filters cover the rest-frame UV to optical spectrum while the JWST filters cover the rest-frame optical to NIR range. As pointed out in Sect. 4.1, the BPASS models are not optimised to reproduce older stars which, we suspect to cause the underestimation of the model fluxes in these specific filters, a result supported by Osborne & Salim (2024). According to this test on our full sample, we conclude that the BC03 models better reproduce the photometric measurements.
We also compare the quality of the fit of the spectroscopic data for both BC03 and BPASS models by comparing the EWs measured on the observed spectra to their bayesian estimation with the CIGALE fit. As can be seen in Fig 8, both models are equally good at reproducing the observed EW larger than Å, fainter emission lines (EW ¡ 10 Å) tend to be slightly underestimated.
To conclude, both stellar models are able to fit well the EWs but photometric data are better fitted with BC03 models.
@@ 5.2 Impact of fitting only photometric data and line emission contributions
When a substantial amount of nebular emission contaminates the broadband data (Anders & Fritze-v. Alvensleben, 2003; Zackrisson et al., 2008; Pacifici et al., 2015), it can lead to large uncertainties on the results of the SED fitting as the code lacks the information on the relative contribution of the line to the broadband emission. Notorious examples of erroneous interpretations due to intense nebular emission are broadband measurements made with Spitzer/IRAC (Fazio et al., 2004) interpreted as a Balmer break (e.g. Chary et al., 2005; Zackrisson et al., 2008; van der Wel et al., 2011) and more recently measurements with JWST which mimic a Lyman break leading to incorrect estimates of redshifts (e.g. Donnan et al., 2023; Arrabal Haro et al., 2023a; Zavala et al., 2023; Davis et al., 2023) . Below, we focus on the consequences of fitting only photometric data which accounts for the vast majority of studies making use of SED fitting. We will also measure the contribution of line emission in some of the NIRCam broadband filters obtained with the fits of both photometric fluxes and EWs. For these studies we fit our 13 galaxies with both BC03 and BPASS models despite the results of the BIC test. We choose to use both as more and more studies include the BPASS models.
@@ 5.2.1 Fitting only photometric data
We start by looking at the quality of the fit of photometric data only. We obtain the same trend as in Fig 7 i.e. an overestimation of the fluxes in the four bluest NIRCam filters with the BPASS models. Next, we compare the estimations of the EWs of hydrogen recombination lines with the measured EWs and we present the results in Fig 9. As expected, these estimations are less accurate than the ones obtained when fitting simultaneously photometric fluxes and EWs (Fig 8). Both stellar populations models tend to overestimate the EWs of the lines: CIGALE is favouring stronger bursts of star formation compared to the ones obtained when fitting the spectro-photometric data: the stellar mass fraction created during the burst, , increases from an average value of 0.05 to 0.13. This overestimation is much more prevalent with the BPASS models. This larger contribution of the lines found for the fits with BPASS (average of 45%) compared to those made with BC03 (average of 15%) is explained by the addition of a more intense and younger burst of star formation with median values of and equal to 0.2 and 50 Myrs for BPASS against 0.1 and 100 Myrs for BC03.
@@ 5.2.2 Contribution of line emission in broad-band filters
To estimate the contribution of the emissions lines to the flux measured in each of our broadband filters, we go back to the fit of photometric data and EWs and we consider the best model obtained with the BC03 and BPASS models. We compute the ratio of the flux coming from the nebular emission and the flux of the continuum within the broadband filter. At the redshift of our galaxies, strong emission lines (H, H, H, [OIII]-4959, [OIII]-5007, [NII]-6548 and [NII]-6583) are lying within the bandpasses of the NIRCam F150W and F200W filters. Out of all these lines, H is the most intense which makes its contribution easier to track. In Table 5 we provide the contribution of the nebular emission 555Other lines than these strong emission lines, as well as the nebular continuum, are included by CIGALE in the nebular emission, but their contribution is found negligible. in the NIRCam broadband filter containing the H line for the BC03 and BPASS models. The filter to which H is contributing changes between F150W and F200W depending on the redshift of each source. We plot the percentage of nebular emission contribution in the filter containing H in Fig 10 against the specific SFRs (sSFRs) of the galaxies. The sSFRs are obtained by dividing the bayesian estimation of the SFRs averaged over the last 10 Myrs with the bayesian estimation of the stellar mass both obtained from the output parameters of the best model from the SED fitting. Four galaxies (IDs 8515, 8736, 16991 and 18294) are not included since the H line is redshifted in-between NIRCam filters.

### v3

@@ 3.2 Analysis of observed line ratios
We want to ascertain if some objects in our sample have hidden attenuation before going further into our analysis. Indeed, a galaxy with a high Pa/H ratio compared to a Balmer decrement close to 2.86 would indicate a hidden and highly attenuated component. To do this, we compare the observed line fluxes ratios to the theoretical ones. However, this comparison comes with a few caveats, the first one being slit losses. Following (Napolitano et al., 2024, submitted to A&A) we assume that the slit losses affect all lines equally. The second issue comes from the underlying absorption lines of the stellar continuum, especially for H, H and H which have an absorption line EW 5Å(see Table 6). We correct the fluxes for these absorptions by computing the absorbed flux from the best model of CIGALE (Sect.3.3).
@@
We find that our objects encounter at most a very moderate hidden attenuation, a result comparable to what Reddy et al. (2023) found for a sample of similar redshift CEERS sources. Taking into account the uncertainties on the flux measurements, our galaxies have either a very low attenuation or are compatible with the attenuation model presented in Sect 3.1. In the left panel of Fig 3, four objects are located to the right of the model lines (IDs 8736, 8515, 10293 and 5409). The H/H of IDs 8736 and 10293 shows that both objects are consistent with the models and their Pa/H ratio is compatible with this result within less than 1 (0.8 and 0.5 respectively). ID 5409 has a large uncertainty on its Pa/H ratio which also makes this galaxy compatible with the models within 1 . Finally, the Pa/H ratio of ID 8515 is compatible with the models within 2 while its other Pa/H ratio, Pa/H, is compatible with the models and a very low amount of attenuation within 0.1 (see the right panel of Fig 3). From Fig 3, we conclude that the modified CF00 recipe is able to reproduce the attenuation processes at work in obscured galaxies.
@@ 3.3 Absorption line EWs
One common difficulty of measuring hydrogen emission line EWs is the presence of an underlying absorption line. In our fitting method, presented in Sect.4.3, we reproduce the EW measurements carried out on observed spectra (Sect.2.1), avoiding any correction for an underlying absorption either on observed or model spectra.
Absorption lines are part of the stellar emission templates used to fit the galaxies and in this section we make use of the best fit model to compute seven hydrogen absorption line EWs for each galaxy of our sample. We use the binary models of BPASS as even the resolution of the high resolution models from Bruzual & Charlot (2003) (BC03) is not enough to accurately model these absorption lines. We show an example of the best model of ID 5430 in Fig 4 and the values of the EWs are reported in Table 6.
@@
The main driver of the strength of absorption line EWs is likely to be the SFH of the galaxy. We find a continuous decrease of the strength of the absorption lines with increasing wavelengths from Pa to H. We explain this result by a larger contribution of old stars to the stellar emission at longer wavelength, reducing the impact of absorption lines from late B and A stars on the stellar continuum.
@@ 4 Spectral energy distribution fitting
To fit the SEDs from the combination of broadband fluxes and EWs, we use the modelling software CIGALE. For a complete description of CIGALE and its functionalities see Boquien et al. (2019). The models are built by successively calling modules, each corresponding to a single physical component or process. CIGALE combines a stellar SED built with an SFH, stellar population models and nebular emission with dust absorption and emission components. The energy balance between stellar and nebular dust absorption and dust re-emission is conserved. Energetic photons produced by massive stars ionise the surrounding gas which re–emits the energy in the form of a series of emission lines and a nebular continuum.
The quality of the fit is assessed by the value of the and the values of the physical parameters and their corresponding uncertainties are estimated as the likelihood-weighted means and standard deviations. Below we briefly present the modules we use and our input parameter values which are summarised in Table 4.
Our SED fitting combines photometric fluxes and EWs as spectroscopic data. When combining integrated photometric data and EWs measured on the NIRSpec spectra we assume that these EWs are representative of the whole galaxies.
Along with the SED fitting process with CIGALE it is possible to build and analyse a mock catalogue related to the observed dataset(Buat et al., 2014; Ciesla et al., 2015). CIGALE uses the best-fit model of each of the objects obtained with the SED-fitting procedure. The flux densities and EWs of the mock SEDs are then computed by randomly picking a flux or EW value from the normal distribution generated using the best model value and the error of the input data as standard deviation. CIGALE is then run on this artificial catalogue with the same configuration as for the first run in order to compare the exact values of the physical parameters corresponding to the artificial SEDs to the parameters estimated by the code with the probability density function of each parameter. Below we use this comparison to check the robustness of the output SED fitting parameters of interest for our study and our ability to constrain them.
@@ 4.1 Stellar population models and star formation history
We consider two models of stellar populations: the widely used BC03 models and the BPASS models (version 2.2) which account for binary stars (Eldridge et al., 2017; Stanway & Eldridge, 2018). The latter has been shown to be very efficient when fitting UV and optical photometry with spectroscopic data (e.g. Larson et al., 2023; Bolamperti et al., 2023; Rezaee et al., 2023; Tacchella et al., 2023; Lecroq et al., 2024) as they were built with a focus on best reproducing the emission from young and massive stars (e.g. Eldridge et al., 2017; Stanway & Eldridge, 2018; Reddy et al., 2022; Marchant & Bodensteiner, 2023, submitted to ARA&A). The BPASS binary models have known issues when fitting stellar populations older than a few gigayears (Eldridge et al., 2017; Stanway & Eldridge, 2018; Han & Han, 2019), the first one being that models remain bluer compared to single star models. Two reasons are given by Eldridge et al. (2017) to explain this effect: the fixed binary fraction regardless of the mass of the star and the width of the time bins at late ages combined with a relatively sparse sampling of stellar masses. Binary models also take a longer time to reach the quiescence (Eldridge et al., 2017). This leads to an underestimation of the number of old and small binary stars.
In both cases, we choose the Initial Mass Function of Chabrier (2003) and we assume that the metallicity is fixed to the solar value: .
@@ 4.2 Dust attenuation and emission
To account for the dust attenuation, we apply the model presented in Sect. 3.1 which corresponds to the modified_CF00 module in CIGALE. In this module, , , and (Eq. 1, 2, and 4) are free parameters. We sample the parameter space of , and very thoroughly (with 0.1 step). ranges from -0.4 to -1.5 to represent the broad distributions of values found from both observations (e.g. Lo Faro et al., 2017; Buat et al., 2019; Pantoni et al., 2021; Seillé et al., 2022) and simulations (e.g. Chevallard et al., 2013; Roebuck et al., 2019; Trayford et al., 2020). The tests we perform on mock catalogues show that is not constrained. We thus introduce only three values for : the widely used value of -1.3 (Charlot & Fall, 2000; da Cunha et al., 2008; Chevallard et al., 2013), -0.7 as proposed in the original CF00 recipe (Charlot & Fall, 2000) and an in-between value of -1.0.
To model the dust re-emission of the five sources that have far-IR data, we use the models from Draine & Li (2007), updated in Draine et al. (2014) which can be found in the dl2014 module in CIGALE. These models reproduce the diffuse dust emission heated by the general stellar population which means that the dust is illuminated with a single radiation field . The models also include a component to recreate the emission of the dust tightly linked to star-forming regions. In that case, the dust is illuminated with a variable radiation field intensity superior to The dust mass fraction of dust linked to the star–forming regions (respectively diffuse emission) is referred to as (respectively 1 - ). The polycyclic aromatic hydrocarbon (PAH) abundance, is also a free parameter in the models.
@@ 4.3 Nebular emission and EWs
CIGALE models the emission of the ionised gas with CLOUDY 17.01 (Ferland et al., 1998, 2017) with the configuration described in Villa-Vélez et al. (2021) and Theulé et al. (2024). 129 nebular emission lines intensities are calculated and re-scaled with the number of Lyman continuum photons from the stellar emission of the galaxy. We are only interested in hydrogen lines which makes the impact of the input parameters on the emission line intensities negligible. Therefore we decide to fix all parameters pertaining to the nebular emission to the values mentioned below. The radiation field intensity is given by the ionisation parameter:
where is the number density of photons capable of ionising hydrogen and the number density of hydrogen. The ionisation parameter is fixed to -2.0, a value typical of extra galactic HII regions (Osterbrock & Ferland, 2006). These configurations are also parameterised according to the gas-phase metallicity, Zgas, which is fixed to the value of the stellar metallicity (Z=0.02). The electron density is set to 100 cm-3, a typical density for the ionised parts of a diffuse nebula (Osterbrock & Ferland, 2006). The ionising spectrum is modelled as a constant burst of star formation of 10 Myrs (Boquien et al., 2019; Villa-Vélez et al., 2021) computed with either BC03 or BPASS models.
We model the Paschen and Balmer line EWs 777The EWs computation used in this work is not included in the current released version of CIGALE, 2022.1. The newest version of the EWs computation is available on request. so as to reproduce the measurements described in Sect 2.1. For each emission line of each galaxy, the input wavelengths (a six values ”array”) defined with LiMe are used to delimit the line region and two nearby featureless continua on each side of the line. The code then performs a linear fit of the overall continuum which will be subtracted from the line. LiMe deblends the H and [NII] doublet of our galaxy sample (see Fig 1) and these lines are also considered separately in CIGALE. Finally as pointed out in Sect 3.2 these EWs are not corrected for underlying absorption as they are computed and thus affected in the same way as the EWs measured with LiMe.
@@ 5.1 Fit of the spectro-photometric data of BC03 vs BPASS
We adopt the limit of ¿ 6 to put a strong preference on a model, a positive preference being defined as 2¡ ¡ 6 (Salmon et al., 2016). The distribution is shown in Fig 6. Eleven objects present a strong preference (and one, ID 10293, a positive preference) for the BC03 models while only one galaxy (ID 23542) is found with a strong preference for the BPASS models. For this object the photometric data is better fitted with the BC03 models (see the SEDs in Appendix A) but the EWs are better fitted with the BPASS models. The uncertainties on the measured EWs are very small for this galaxy ( times lower than for the other sources) which leads to a higher impact of spectroscopic data on the calculation of the value.
We present a more detailed comparison of the quality of the fit of the photometric data with BPASS and BC03 models in Fig 7. For each filter and both models we compute the ratio between the bayesian estimate of the flux given by CIGALE over the measured flux. While the CFHT and HST fluxes do not show a significant difference of quality between both models, there is a clear trend for the four bluest NIRCam filters to underestimate the model fluxes with respect to the observed values when the binary BPASS models are used. The CFHT and HST filters cover the rest-frame UV to optical spectrum while the JWST filters cover the rest-frame optical to NIR range. As pointed out in Sect. 4.1, the BPASS models are not optimised to reproduce older stars which we suspect to cause the underestimation of the model fluxes in these specific filters, a result supported by Osborne & Salim (2024). According to this test on our full sample we conclude that the BC03 models better reproduce the photometric measurements.
We also compare the quality of the fit of the spectroscopic data for both BC03 and BPASS models by comparing the EWs measured on the observed spectra to their bayesian estimation with the CIGALE fit. As can be seen in Fig 8 both models are equally good at reproducing the observed EW larger than Åwhile fainter emission lines (EW ¡ 10 Å) tend to be slightly underestimated.
To conclude both stellar models are able to fit well the EWs but photometric data are better fitted with BC03 models.
@@ 5.2 Impact of fitting only photometric data and line emission contributions
When a substantial amount of nebular emission contaminates the broadband data (Anders & Fritze-v. Alvensleben, 2003; Zackrisson et al., 2008; Pacifici et al., 2015), it can lead to large uncertainties on the results of the SED fitting as the code lacks the information on the relative contribution of the line to the broadband emission. Notorious examples of erroneous interpretations due to intense nebular emission are broadband measurements made with Spitzer/IRAC (Fazio et al., 2004) interpreted as a Balmer break (e.g. Chary et al., 2005; Zackrisson et al., 2008; van der Wel et al., 2011) and more recently measurements with JWST which mimic a Lyman break leading to incorrect estimates of redshifts (e.g. Donnan et al., 2023; Arrabal Haro et al., 2023a; Zavala et al., 2023; Davis et al., 2023, submitted to ApJ) . Below we focus on the consequences of fitting only photometric data which accounts for the vast majority of studies making use of SED fitting. We also measure the contribution of line emission in some of the NIRCam broadband filters obtained with the fits of both photometric fluxes and EWs. For these studies we fit our 13 galaxies with both BC03 and BPASS models despite the results of the BIC test. We choose to use both as more and more studies include the BPASS models.
@@ 5.2.1 Fitting only photometric data
We start by looking at the quality of the fit of photometric data only. We obtain the same trend as in Fig 7 i.e. an overestimation of the fluxes in the four bluest NIRCam filters with the BPASS models. Next we compare the estimations of the EWs of hydrogen recombination lines with the measured EWs and we present the results in Fig 9. As expected these estimations are less accurate than the ones obtained when fitting simultaneously photometric fluxes and EWs (Fig 8). Both stellar population models tend to overestimate the EWs of the lines as CIGALE is favouring stronger bursts of star formation compared to the ones obtained when fitting the spectro-photometric data: the stellar mass fraction created during the burst increases from an average value of 0.05 to 0.13. This overestimation is much more prevalent with the BPASS models. This larger contribution of the lines we find for the fits with BPASS (average of 45%) compared to those made with BC03 (average of 15%) is explained by the addition of a more intense and younger burst of star formation with median values of and equal to 0.2 and 50 Myrs for BPASS against 0.1 and 100 Myrs for BC03.
@@ 5.2.2 Contribution of line emission in broad-band filters
To estimate the contribution of the emissions lines to the flux measured in each of our broadband filters, we go back to the fit of photometric data and EWs and we consider the best model obtained with the BC03 and BPASS models. We compute the ratio of the flux coming from the nebular emission and the flux of the continuum within the broadband filter. At the redshift of our galaxies, strong emission lines (H, H, H, [OIII]-4959, [OIII]-5007, [NII]-6548 and [NII]-6583) are lying within the bandpasses of the NIRCam F150W and F200W filters. Out of all these lines, H is the most intense which makes its contribution easier to track. In Table 8 we provide the contribution of the nebular emission 999Other lines than these strong emission lines, as well as the nebular continuum, are included by CIGALE in the nebular emission, but their contribution is found negligible. in the NIRCam broadband filter containing the H line for the BC03 and BPASS models. The filter to which H is contributing changes between F150W and F200W depending on the redshift of each source. We plot the percentage of nebular emission contribution in the filter containing H in Fig 10 against the specific SFRs (sSFRs) of the galaxies. The sSFRs are obtained by dividing the bayesian estimation of the SFRs averaged over the last 10 Myrs with the bayesian estimation of the stellar mass both obtained from the output parameters of the best model from the SED fitting. Four galaxies (IDs 8515, 8736, 16991 and 18294) are not included since the H line is redshifted in-between NIRCam filters.